OpenStudy (anonymous):

F(X)= integral from sqrtx to 2 x Tan^-1 t dt... Please example steps thanks

OpenStudy (anonymous):

\[F(x)=\int_{\sqrt{x}}^{2x} \arctan (t) \ \text{d}t\]

OpenStudy (anonymous):

what do u want to do with this function?

OpenStudy (anonymous):

Im trying to solve the problem

OpenStudy (anonymous):

im not sure the first step of this equation

OpenStudy (anonymous):

use integration by parts

OpenStudy (anonymous):

such as...

OpenStudy (anonymous):

i completely forgot Calcu i took it years ago and im trying to remember

OpenStudy (anonymous):

1 . arctan(t)......assume u= arctan(t) and v=1

OpenStudy (anonymous):

wait where v=1 come from?

OpenStudy (anonymous):

is v the derivates of arctant?

OpenStudy (anonymous):

\[\int\limits_{a}^{b} uv dx = u \int\limits vdx - \int\limits (du/dx) \int\limits (vdx) dx\]

OpenStudy (anonymous):

here u= arctan(t) and v=1

OpenStudy (anonymous):

it makes sense to u?

OpenStudy (anonymous):

im not understanding where v=1 came from i understand the formula but unsure of v=1

OpenStudy (anonymous):

integrand was arctan(t)...if we multiply it by one our integrand wont change

OpenStudy (anonymous):

we multiplied the integrand by one for using the integration by parts...

OpenStudy (anonymous):

ok...i think i understand

OpenStudy (anonymous):

and after integration...put the limits

OpenStudy (anonymous):

so this would work like chain rule

OpenStudy (anonymous):

yes u can say that...

OpenStudy (anonymous):

ok but im trying to compute the equations into the steps...

OpenStudy (anonymous):

look for Bernoulli's theorem...

OpenStudy (anonymous):

ok